OpenStudy (anonymous):

Write the linear equation y+3=2/5(x +9) in Standard Form.

OpenStudy (anonymous):

5x+45+y+3-2=0 5x+y+46=0

OpenStudy (anonymous):

thats not the answer @Muskan

OpenStudy (lgbasallote):

\[y +3 = \frac 25(x+9)\] is that the questioN?

OpenStudy (anonymous):

Yes @lgbasallote

OpenStudy (lgbasallote):

multiply the equation by 5 what do you get?

OpenStudy (anonymous):

5y+3=2x+18 -2x+5y-15=0 is this rite?@Dee_123

OpenStudy (anonymous):

Im Trying To Get It In Standard Form

OpenStudy (lgbasallote):

i know

OpenStudy (lgbasallote):

do what i said

OpenStudy (lgbasallote):

what do you get?

OpenStudy (anonymous):

5y+3=2x+18 -2x+5y-15=0 is this rite? @Dee_123

OpenStudy (ash2326):

@Dee_123 did you get it?